# Holographic Integrated Data and Energy Transfer

### v1

@@ IV-B Multi-user H-IDET system
(a) The normalized beam pattern of the whole H-MIMO
(b) The normalized beam pattern of the beamforming for EUs
(c) The normalized beam pattern of the beamforming for DU1
(d) The normalized beam pattern of the beamforming for DU2
Figure 9: Near-filed focusing for the IDET system
We then analyse the nearfield focus effect for the H-IDET system in Fig. 9. Generally, the nearfield range is related to the equivalent array width. Without loss of generality, we set $L_{x}=1.5\mathrm{m},L_{y}=0.03\mathrm{m}$ to reduce the computational complexity of simulation. Two DUs are located at (±$(\pm 5,0,25)$ while two EUs are located at (±$(\pm 1,0,1)$. The received power at each location is normalized by the path loss. Observe from Fig. 9(a) that the integrated H-IDET focus energy at the position of EUs and DUs respectively. Specifically, the beam pattern of EUs is illustrated in Fig. 9(b), where the beamforming of EUs only focuses beams at the EUs and achieves the spatial filtering effect for the DUs. Moreover, the beam pattern of DUs are illustrated in Fig. 9(c) and Fig. 9(d), respectively, where the beams are focused at their target DU. At the same time, the beams also points to the EUs and achieve the spatial filtering effect for other DUs.

### v2

@@ IV-B Multi-user H-IDET system
We then analyse the near-field focus effect for the H-IDET system in Fig. 9. Generally, the near-field range is related to the equivalent array width. Without loss of generality, we set $L_{x}=1.5\mathrm{m},L_{y}=0.03\mathrm{m}$ to reduce the computational complexity of simulation. Two DUs are located at (±$(\pm 5,0,25)$ while two EUs are located at (±$(\pm 1,0,1)$. The received power at each location is normalized by the path loss. Observe from Fig. 9(a) that the integrated H-IDET is able to focus energy at the position of EUs and DUs respectively. Specifically, the beam pattern of EUs is illustrated in Fig. 9(b), where the beamforming of EUs only focuses beams at the EUs and achieves the spatial filtering effect for the DUs. Moreover, the beam pattern of DUs are illustrated in Fig. 9(c) and Fig. 9(d), respectively, where the beams are focused at their target DU. At the same time, the beams also point to the EUs for better WET performance and achieve the spatial filtering effect for other DUs.